Compute the present value of a perpetuity of $16,000 per year given an interest rate of 4 percent p.a., compounded monthly.
Answer #1

Effective annual rate = ( 1+ 4%/12)12-1 = 4.07415%

PV of perpetuity = 16,000/ ( 1+ 4%/12)12-1 = 392,719.54

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well